Two machines are turned on at 8:00 am If one can produce 45 items each hour and the other can produce 71 items each hour, at what time will they produce a total of 232 items?
Answer by rfer(16322)

find the average per hour divided into the total number of items gives the number of hours. Add that to 8am for compltion time.
71+45=116/2=58
232/58=4hrs or noon